# The formation of Schrödinger cat-like states in the process of spontaneous parametric down-conversion

### v5

@@ Abstract
We investigate the formation of entangled Schrödinger cat-like states (SCLSs) in both the fundamental and second-harmonic modes during spontaneous parametric down-conversion (SPDC), considering quantum depletion of the pump. Unlike conventional semiclassical treatments that assume an undepleted classical pump, we fully quantize both modes and explore their dynamics under non-dissipative and dissipative regimes. Using the Wigner function, we visualize the non-Gaussian character and interference patterns indicative of these entangled states. Our analysis shows significant squeezing (up to $3.76\,\mathrm{dB}$ without dissipation and $3.65\,\mathrm{dB}$ with dissipation) and super-Poissonian statistics (Fano factor $\gg 1$), both persisting under moderate losses—features that extend beyond those of conventional optical Schrödinger cat states, which lack such robustness. In contrast to semiclassical models yielding statistical mixtures, our approach captures energy exchange between modes, phase-dependent evolution, interference, and interaction-length sensitivity. These features are critical for continuous-variable quantum technologies. The generated entangled SCLSs are highly relevant for applications in quantum sensing and quantum information processing. Our findings pave the way for realizing strongly non-Gaussian entangled states using $\chi^{(2)}$ media at moderate pump intensities, offering advantages over $\chi^{(3)}$-based schemes.
preprint: APS/123-QED
I Introduction
@@
We investigate the formation of entangled Schrödinger cat-like states (SCLSs) in the fundamental and second-harmonic modes of spontaneous parametric down-conversion (SPDC) under a fully quantized pump model. Unlike conventional optical Schrödinger cat states (even/odd coherent states [Dodonov1974]) that exhibit only non-classical interference, our SCLSs demonstrate additional quantum features including significant squeezing (up to $3.76\,\mathrm{dB}$) and super-Poissonian statistics (Fano factor $\gg 1$), both robust under moderate losses. These properties, combined with their inherent entanglement between modes, distinguish them fundamentally from single-mode cat states. Going beyond undepleted-pump approximations, we analyze their dynamics in non-dissipative and dissipative regimes. Wigner function visualizations reveal the non-Gaussian character and interference patterns of these states, confirming their SCLS nature. Unlike semiclassical models that produce statistical mixtures, our approach captures energy exchange, phase-dependent evolution, and interference, highlighting critical properties for continuous-variable quantum technologies. These entangled SCLSs are promising for quantum sensing and information processing, offering a path to strong non-Gaussianity in $\chi^{(2)}$ media at moderate pump intensities.
@@
While the standard treatment of SPDC approximates the Hamiltonian as quadratic by treating the pump classically, our work employs the full cubic Hamiltonian $\hat{H}_{\mathrm{int}}\propto\hat{a}_{1}^{\dagger 2}\hat{a}_{2}+\mathrm{h.c.}$ with a quantized pump. The key novelty lies in exploring the high-depletion operating condition—where specific initial states lead to approximately 64% pump conversion—whose quantum features had remained largely uncharacterized due to the dominance of the undepleted-pump approximation. This operating condition could be realized with modern highly-efficient nonlinear crystals (e.g., periodically poled structures [see Belinsky2018, and references therein]) and using contemporary numerical tools like QuTiP [johansson2012qutip, JOHANSSON20131234].
@@
Schrödinger cat-like states, characterized by negative values of the Wigner function, play an important role in, e.g., observing phase displacements due to their phase sensitivity [Shukla2023, Toscano2006, Salykina2023, Singh2024]. Schrödinger cat states can be used to encode cat qubits [Cochrane1999] and building Ising machines [Yamamoto2020].
@@
In the field of quantum nonlinear optics, the generation of non-Gaussian states requires at least the cubic form of the interaction Hamiltonian [Braunstein2005]. The formation of non-Gaussian states [Sychev2017] (e.g., Schrödinger cat-like states) has been considered in [Nikitin1991, Singh2021] during quantum state evolution in second harmonic generation with $\chi^{(2)}$ media, with [Singh2021] demonstrating interference patterns using Wigner functions.
@@
The classical and non-depleted or semi-classical method approximates the interaction Hamiltonian for the parametric down-conversion (PDC) process, which is based on $\chi^{(2)}$ (and is cubic), by a quadratic one (i.e., pump mode is considered classic and undepleted). Such an approximation neglects the non-Gaussianity features present in the initial cubic form of the interaction Hamiltonian. Another approximation method, e.g., the expansion of unitary operators [Agarwal1974, Beskrovnyi1996, Belinsky2018], can be applied to consider a full treatment of quantum effects present in all the modes of the PDC process, but starts to deviate at later stages [Belinsky2018]. However, the diagonalization method [Walls1970, Walls1972, Gantsog1991, Tanas1992, Belinsky2018] can be used to study evolution at longer interaction lengths without losing the quantum effects present in the interacting modes.
@@
Previous studies have implemented a full quantum mechanical treatment of parametric down-conversion (PDC) with a depleted pump mode [Walls1970, Walls1972, Gantsog1991, Tanas1991, Tanas1992, Drobny1992], investigating statistical properties including mean photon numbers, quadrature fluctuations in the fundamental mode, and phase properties of both modes [Gantsog1993].
@@
Conversely, semiclassical treatments [w1988, Walls2008] assume a classical, undepleted pump, leading to a Fokker-Planck equation whose steady-state solution yields a statistical mixture of two coherent states. This mixture inherently lacks crucial quantum features: it exhibits no interference pattern, cannot form genuine even or odd cat states, omits energy exchange between modes, shows no dependence on interaction length, and is confined to the dissipative regime.
@@
In contrast, our work demonstrates the formation of SCLSs in both modes across both non-dissipative and dissipative regimes. Our results demonstrate the preservation of the quantum interference structure and provide clear evidence of energy exchange between the fundamental and second-harmonic modes (see Fig. 1 and Fig. 2).
@@
In this paper, we have studied the formation of SCLSs for two cases (non-dissipative and dissipative) when both modes (fundamental and second harmonic) are considered as quantum and depleted during the effective realization of the SPDC process based on $\chi^{(2)}$ [Walls2008, Agarwal2013]. QuTiP [johansson2012qutip, JOHANSSON20131234] is used to numerically solve the Lindbladian superoperator responsible for the SPDC process. The formation of SCLSs is studied and illustrated qualitatively using the Wigner function. Quantitative analysis of the SCLSs is performed by calculating squeezing levels (variances of quadrature components and Fano factor), values of photon number distributions and the entanglement between the two modes. The SCLSs are observed in the fundamental and second harmonic modes for non-dissipative and dissipative cases when the fundamental mode is in the vacuum state and the second harmonic mode is in the coherent state.
II SPDC under pump depletion and losses
@@
Let three stationary degenerate monochromatic optical plane wave modes $\hat{a}_{s},\hat{a}_{i},\hat{a}_{p}$ of frequencies $\omega_{s}$, $\omega_{i}$, $\omega_{p}$ propagate collinearly in a nonlinear optical crystal with non-zero susceptibility $\chi^{(2)}$. Subscripts $i,s,p$ belong to the idler, signal, and pump modes. In case of degenerate frequencies: $\omega_{s}=\omega_{i}=\omega$, $\omega_{p}=2\omega$, $\hat{a}_{i}=\hat{a}_{s}=\hat{a}_{1},\hat{a}_{p}=\hat{a}_{2}$. For effective realization of the SPDC process, it is assumed that all three modes can be phase-matched [Dmitriev1999] or quasi-phase-matched [Belinsky2018]. The interaction Hamiltonian of the SPDC process $(2\omega=\omega+\omega)$ [Walls2008, Agarwal2013] is given by
$\hat{H}_{int}=\hbar g(\hat{a}^{2}_{1}\hat{a}_{2}^{\dagger}+h.c.),$ (1)
@@
The Lindblad master equation for the density matrix $\hat{\rho}$ describing an open quantum system is used to study the quantum dynamics of the interaction Hamiltonian (1 and can be written as [Agarwal2013, Walls2008, BP2002, Strinati2024]
@@
where $\tau=gt$ is the normalized interaction length. $\hat{C}_{j}=\sqrt{\gamma_{j}/g}\hat{a}_{i}$ are the Lindblad operators describing the dissipative part of the dynamics, where $\gamma_{j}\geq 0$ are the cavity damping rates of the modes and $g>0$.
@@
III Formation of cat-like states in both modes
One can study and analyze the quantum statistical properties of the modes by using the Wigner quasiprobability distribution [Walls2008, Agarwal2013]. Phase space portraits of the Wigner function help to visualize delicate patterns, such as (a) interference (wave nature) present in the superposition of macroscopically distinct states, e.g., SCLSs, (b) non-Gaussianity of the state, i.e., the negative value of the Wigner function. The Wigner functions of the modes $\hat{a}_{1}$ and $\hat{a}_{2}$ are calculated using
@@ III Formation of cat-like states in both modes
Figure 1: Wigner functions $W_{1}(\alpha_{1},\tau)$ and $W_{2}(\alpha_{2},\tau)$ during SPDC at $\tau=0.38$. Top row (a,b): Non-dissipative ($\gamma_{1,2}=0$), $N_{1}\approx 25.2$, $N_{2}\approx 7.36$. Bottom row (c,d): Dissipative ($\gamma_{1,2}=0.15$), $N_{1}\approx 23.88$, $N_{2}\approx 6.95$. Initial: $\hat{a}_{1}$ vacuum, $\hat{a}_{2}$ coherent state ($|\alpha_{20}|^{2}=20$, $\varphi_{20}=\pi/2$).
@@
IV Application-relevant statistical properties
@@
IV.1 Mean number extrema as indicators of SCLSs formation. Squeezing in both modes
@@
where $\hat{x}_{j}=2^{-(1/2)}(\hat{a}_{j}+\hat{a}^{\dagger}_{j})$ and $\hat{p}_{j}=-i2^{-(1/2)}(\hat{a}_{j}-\hat{a}^{\dagger}_{j})$ are the quadrature components of the modes $\hat{a}_{j}$. Fig. 2 shows the evolution of both (i) the mean photon numbers (Eq. 4) and (ii) the quadrature component variances (Eqs. 5 and 6) for the non-dissipative case. SCLSs form at the first extremal values of the photon numbers: the maximum for mode $\hat{a}_{1}$ and the minimum for mode $\hat{a}_{2}$, both occurring at $\tau=0.38$. Notably, the quadrature variances $\Delta^{2}p_{1}(0.38)=0.23$ (mode $\hat{a}_{1}$) and $\Delta^{2}x_{2}(0.38)=0.21$ (mode $\hat{a}_{2}$) demonstrate squeezing, corresponding to 3.37-3.37 dB and 3.76-3.76 dB respectively.
Figure 2: Evolution of (a) $N_{j}$ and (b) quadrature variances SPDC process. $N_{1}$ ($\hat{a}_{1}$, solid black, -) peaks while $N_{2}$ ($\hat{a}_{2}$, dashed black, --) minimizes at $\tau=0.38$. Variances: $\Delta^{2}x_{1}$ (blue $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{1}$ (blue $\cdot$-), $\Delta^{2}x_{2}$ (red $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{2}$ (red $\cdot$-). $\bigstar$ marks extrema.
Figure 3: Photon number distributions $P_{j}(n_{j})$ at $\tau=0.38$: (a,b) non-dissipative ($\gamma_{1,2}=0$) and (c,d) dissipative ($\gamma_{1,2}=0.15$) regimes for modes $\hat{a}_{1}$ (a,c) and $\hat{a}_{2}$ (b,d).
IV.2 Dissipation-induced parity mixing in SCLSs
@@
IV.3 Robustness of super-Poissonian statistics to dissipation
The value of the Fano factor can identify the type (sub-Poissonian, Poissonian, super-Poissonian) of the distributions of the studied SCLS $\hat{\rho}_{j}$. Thus, the Fano factor [Vogel] is calculated using

### v6

# The formation of entangled Schrödinger cat-like states in the process of spontaneous parametric down-conversion

@@ Abstract
We investigate entangled Schrödinger cat-like states (SCLSs) in degenerate and non-degenerate spontaneous parametric down-conversion (SPDC) with a fully quantized, depleted pump. Our fully quantum treatment, visualized via Wigner functions, reveals non-Gaussian features and interference patterns absent in semiclassical models. For degenerate SPDC, we demonstrate significant squeezing (up to $4.04\,\mathrm{dB}$) and robust super-Poissonian statistics in both non-dissipative and dissipative regimes. Extending to non-degenerate SPDC, we confirm that pump quantization also generates non-Gaussian states in all modes and yields a higher-dimensional entanglement structure, evidenced by a larger Schmidt number ($K^{(\mathrm{ND})}\approx 10.38$) compared to the degenerate case ($K\approx 1.93$). Our approach captures critical dynamics like energy exchange and phase-dependent evolution. These entangled SCLSs, non-Gaussian states realizable in $\chi^{(2)}$ media at moderate intensities and offering advantages over $\chi^{(3)}$-based schemes, are promising resources for quantum sensing and information processing.
Keywords: Schrödinger cat states; SPDC (degenerate & non-degenerate); quantum pump depletion; entanglement; non-Gaussian states; quadrature squeezing
1 Introduction
We investigate the formation of entangled Schrödinger cat-like states (SCLSs), non-Gaussian states in the fundamental and second-harmonic modes of spontaneous parametric down-conversion (SPDC) under a fully quantized pump model. Unlike conventional optical Schrödinger cat states (even/odd coherent states [1]) that exhibit only non-classical interference, our SCLSs demonstrate additional quantum features including significant squeezing (up to $4.04\,\mathrm{dB}$) and super-Poissonian statistics (Fano factor $\gg 1$), both robust under moderate losses. These properties, combined with their inherent entanglement between modes, distinguish them fundamentally from single-mode cat states. Going beyond undepleted-pump approximations, we analyze their dynamics in non-dissipative and dissipative regimes. Wigner function visualizations reveal the non-Gaussian character and interference patterns of these states, confirming their SCLS nature. Unlike semiclassical models that produce statistical mixtures [2, 3], our approach captures energy exchange, phase-dependent evolution, and interference, highlighting critical properties for continuous-variable quantum technologies [4].
These entangled SCLSs, non-Gaussian states are promising for quantum sensing and information processing, offering a path to strong non-Gaussianity in $\chi^{(2)}$ media at moderate pump intensities. Furthermore, we extend our investigation to the non-degenerate (NDSPDC) configuration, demonstrating that pump quantization similarly generates non-Gaussian resources and yields richer, tripartite entanglement structures compared to the degenerate case.
@@ 1 Introduction
While the standard treatment of SPDC approximates the Hamiltonian as quadratic by treating the pump classically, our work employs the full cubic Hamiltonian $\hat{H}_{\mathrm{int}}\propto\hat{a}_{1}^{2}\hat{a}_{2}^{\dagger}+\mathrm{h.c.}$ or $\hat{H}_{\mathrm{int}}^{(\mathrm{ND})}\propto\hat{a}_{s}\hat{a}_{i}\hat{a}_{p}^{\dagger}+\mathrm{h.c.}$ with a quantized pump. The key novelty lies in exploring the high-depletion operating condition—where specific initial states lead to approximately 63.2% pump conversion (for degenerate SPDC process) — whose quantum features had remained largely uncharacterized due to the dominance of the undepleted-pump approximation. This operating condition could be realized with modern highly-efficient nonlinear crystals (e.g., periodically poled structures [see 5, and references therein]) and using contemporary numerical tools like QuTiP [6, 7].
@@
SCLSs, characterized by negative values of the Wigner function, play an important role in, e.g., observing phase displacements due to their phase sensitivity [8, 9, 10, 11]. Schrödinger cat states can be used to encode cat qubits [12] and building Ising machines [13].
@@
In the field of quantum nonlinear optics, the generation of non-Gaussian states requires at least the cubic form of the interaction Hamiltonian [4]. The formation of non-Gaussian states [14] (e.g., SCLSs) has been considered in [15, 16] during quantum state evolution in second harmonic generation with $\chi^{(2)}$ media, with [16] demonstrating interference patterns using Wigner functions.
@@
The classical and non-depleted or semi-classical method approximates the interaction Hamiltonian for the parametric down-conversion (PDC) process, which is based on $\chi^{(2)}$ (and is cubic), by a quadratic one (i.e., pump mode is considered classic and undepleted). Such an approximation neglects the non-Gaussianity features present in the initial cubic form of the interaction Hamiltonian. Another approximation method, e.g., the expansion of unitary operators [17, 18, 5], can be applied to consider a full treatment of quantum effects present in all the modes of the PDC process, but starts to deviate at later stages [5]. However, the diagonalization method [19, 20, 21, 22, 5] can be used to study evolution at longer interaction lengths without losing the quantum effects present in the interacting modes.
@@
Previous studies implemented fully quantum mechanical treatments of parametric down-conversion (PDC) with pump depletion, explicitly accounting for the correlated exchange of quanta between the pump and down-converted modes [19, 20, 21, 23, 22, 24]. These works investigated statistical properties such as mean photon numbers, quadrature fluctuations in the fundamental mode, and phase properties of both the pump and signal fields [25].
Subsequent theoretical analyses clarified the limitations of the parametric approximation by demonstrating that its validity is governed not by the absence of pump depletion, but by the preservation of pump coherence after the interaction [26]. Extensions of this insight to experimentally relevant regimes examined the role of quantum fluctuations, non-Gaussian state generation, and critical behavior near threshold in optical parametric oscillators [27]. More recent studies have further explored pump depletion, entanglement, and squeezing dynamics beyond the parametric approximation in macroscopic down-conversion processes, identifying the characteristic timescales over which genuinely quantum features emerge simultaneously in the pump and down-converted modes [28].
@@
Semiclassical treatments of PDC [2, 3] assume a classical, undepleted pump field, leading to a Fokker–Planck description whose steady-state solution corresponds to a statistical mixture of two coherent states. While such approaches capture mean-field amplification and phase bistability, this description inherently misses essential quantum features: it exhibits no interference structure, cannot represent genuine even or odd Schrödinger-cat states, neglects energy exchange and back-action between modes, shows no dependence on interaction length, and remains confined to a dissipative steady-state regime.
@@
Our full quantum treatment overcomes these limitations, demonstrating the formation of SCLSs in both modes across both non-dissipative and dissipative regimes. Crucially, our results preserve the quantum interference structure and provide direct evidence of energy exchange between the fundamental and second-harmonic modes (see Fig. 1 and Fig. 2).
@@
In this paper, we have studied the formation of entangled SCLSs for two cases (non-dissipative and dissipative) when both modes (fundamental and second harmonic) are considered as quantum and depleted during the effective realization of the degenerate SPDC process based on $\chi^{(2)}$ [3, 29]. We extend this methodology to the non-degenerate SPDC (NDSPDC) configuration, analyzing the formation of non-Gaussian states and entanglement in its tripartite system under the same fully quantum framework. QuTiP [6, 7] is used to numerically solve the Lindbladian superoperator responsible for the SPDC process. The formation of SCLSs is studied and illustrated qualitatively using the Wigner function. For the degenerate SPDC process, quantitative analysis of the SCLSs is performed by calculating squeezing levels (variances of quadrature components and Fano factor), values of photon number distributions, and the entanglement between the two modes. The SCLSs are observed in the fundamental and second harmonic modes for non-dissipative and dissipative cases when the fundamental mode is in the vacuum state and the second harmonic mode is in the coherent state. For the NDSPDC process, we similarly compute Wigner functions and quantify entanglement via Schmidt number analysis. In this configuration, we observe the formation of non-Gaussian states: the signal mode exhibits a non-Gaussian structure distinct from a thermal state, while the pump mode develops SCLS features.
2 Degenerate SPDC under pump depletion and losses
@@
Let three stationary degenerate monochromatic optical plane wave modes $\hat{a}_{s},\hat{a}_{i},\hat{a}_{p}$ of frequencies $\omega_{s}$, $\omega_{i}$, $\omega_{p}$ propagate collinearly in a nonlinear optical crystal with non-zero susceptibility $\chi^{(2)}$. Subscripts $i,s,p$ belong to the idler, signal, and pump modes. In case of degenerate frequencies: $\omega_{s}=\omega_{i}=\omega$, $\omega_{p}=2\omega$, $\hat{a}_{i}=\hat{a}_{s}=\hat{a}_{1},\hat{a}_{p}=\hat{a}_{2}$. For effective realization of the SPDC process, it is assumed that all three modes can be phase-matched [30] or quasi-phase-matched [5]. The interaction Hamiltonian for the degenerate SPDC process $(2\omega=\omega+\omega)$ [26, 3, 29]
$\hat{H}_{int}=\hbar g(\hat{a}^{2}_{1}\hat{a}_{2}^{\dagger}+\mathrm{h.c.}),$ (1)
@@
The Lindblad master equation for the density matrix $\hat{\rho}$ describing an open quantum system is used to study the quantum dynamics of the interaction Hamiltonian (1) and can be written as [29, 3, 31, 32]
@@
where $\tau=gt$ is the normalized interaction length. $\hat{C}_{j}=\sqrt{\gamma_{j}/g}\hat{a}_{j}$ are the Lindblad operators describing the dissipative part of the dynamics, where $\gamma_{j}\geq 0$ are the cavity damping rates of the modes and $g>0$.
@@
3 Formation of cat-like states in both modes
One can study and analyze the quantum statistical properties of the modes by using the Wigner quasiprobability distribution [3, 29]. Phase space portraits of the Wigner function help to visualize delicate patterns, such as (a) interference (wave nature) present in the superposition of macroscopically distinct states, e.g., SCLSs, (b) non-Gaussianity of the state, i.e., the negative value of the Wigner function. The Wigner functions of the modes $\hat{a}_{1}$ and $\hat{a}_{2}$ are calculated using
@@ 3 Formation of cat-like states in both modes
Figure 1: Wigner functions $W_{1}(\alpha_{1},\tau)$ and $W_{2}(\alpha_{2},\tau)$ for degenerate SPDC process at $\tau=0.38$. Top row (a,b): Non-dissipative ($\gamma_{1,2}=0$), $N_{1}\approx 25.28$, $N_{2}\approx 7.35$. Bottom row (c,d): Dissipative ($\gamma_{1,2}=0.15$), $N_{1}\approx 23.88$, $N_{2}\approx 6.94$. Initial: $\hat{a}_{1}$ vacuum, $\hat{a}_{2}$ coherent state ($|\alpha_{20}|^{2}=20$, $\varphi_{20}=\pi/2$).
@@
4 Application-relevant statistical properties
@@
4.1 Mean number extrema as indicators of SCLSs formation. Squeezing in both modes
@@
where $\hat{x}_{j}=2^{-(1/2)}(\hat{a}_{j}+\hat{a}^{\dagger}_{j})$ and $\hat{p}_{j}=-i2^{-(1/2)}(\hat{a}_{j}-\hat{a}^{\dagger}_{j})$ are the quadrature components of the modes $\hat{a}_{j}$. Fig. 2 shows the evolution of both (i) the mean photon numbers (Eq. 4) and (ii) the quadrature component variances (Eqs. 5 and 6) for the non-dissipative case. SCLSs form at the first extremal values of the photon numbers: the maximum for mode $\hat{a}_{1}$ and the minimum for mode $\hat{a}_{2}$, both occurring at $\tau=0.38$. Notably, the quadrature variances $\Delta^{2}p_{1}(0.38)\approx 0.1976$ (mode $\hat{a}_{1}$) and $\Delta^{2}x_{2}(0.38)\approx 0.1970$ (mode $\hat{a}_{2}$) demonstrate squeezing, corresponding to $4.03$ dB and $4.04$ dB respectively.
Figure 2: Evolution of (a) $N_{j}$ and (b) quadrature variances for degenerate SPDC process. $N_{1}$ ($\hat{a}_{1}$, solid black, -) peaks while $N_{2}$ ($\hat{a}_{2}$, dashed black, --) minimizes at $\tau=0.38$. Variances: $\Delta^{2}x_{1}$ (blue $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{1}$ (blue $\cdot$-), $\Delta^{2}x_{2}$ (red $\cdot{\cdot}{\cdot}$), $\Delta^{2}p_{2}$ (red $\cdot$-). $\bigstar$ marks extrema.
Figure 3: Photon number distributions $P_{j}(n_{j})$ at $\tau=0.38$ for degenerate SPDC process: (a,b) non-dissipative ($\gamma_{1,2}=0$) and (c,d) dissipative ($\gamma_{1,2}=0.15$) regimes for modes $\hat{a}_{1}$ (a,c) and $\hat{a}_{2}$ (b,d).
4.2 Dissipation-induced parity mixing in SCLSs
@@
4.3 Robustness of super-Poissonian statistics to dissipation
The value of the Fano factor can identify the type (sub-Poissonian, Poissonian, super-Poissonian) of the distributions of the studied SCLS $\hat{\rho}_{j}$. Thus, the Fano factor [33] is calculated using